# Semblance study of force-field and convection-diffusion solutions with observations of solar magnetic phenomena

### v2

@@ Abstract
A quantitative analytical model based on the semblance method between the modulation factor with solar phenomena is proposed. Different Local Interstellar Spectra (LIS) have been computed to introduce into a transport equation solution which in turn have been introduced into the atmospheric yield function (Caballero-Lopez & Moraal, 2012), that allows to compute a Cosmic Rays (CR) Modulation Factor. The results were as expected; there are correlation between modulation factor and sunspots, and anticorrelation between modulation factor and mean magnetic field.
A transport equation’s solution is necessary to compute atmospheric yield function, in this case the used transport equation’s solutions were convection-diffusion and force field. Both solutions offer similar models, yet the force field solution shows a higher correlation value in the semblance than the convection-diffusion solution.
Several LIS were also computed because they are introduced into the transport equation solutions. The used LIS were Lagner, Potgieter and Webber LIS in 2003, Burguer and Potgieter LIS in 2000, Garcia-Munoz, Mason and Simpson LIS in 1975 and Ghelfi, Barao, Derome and Maurin LIS in 2017. Those LIS were used because they have a model for different nuclear composition: Helium and Hydrogen. The LIS with more changes when is introduced into the semblance is Garcia-Munoz, Mason and Simpson in 1975.
The atmospheric yield function also need data from some neutron detectors that have changes according their cut-off rigidity value, hence two analyses were made. In the first analysis were selected three stations from different locations: close from the Arctic Circle as OULU station, in the Antarctic as SNAE and close of the equator as AATB station. In the second analysis, five different stations were averaged by arithmetic mean to observe and analyze what is the general behaviour of arrival cosmic rays on the Earth. The computed semblance models are very similar with significant differences that are shown in an analysis of zeros.
Many images were produced when the analyses were computed, then the main images, that are used as examples, are shown along the paper, while the rest can be found in the appendix.
Cosmic Rays — Convetion-Diffusion solution — Force field solution — Semblance — Geophysics — Local interstellar spectal — Atmospheric yiel funtion
@@ 1 Introduction
Cosmic rays are electrically charged particles capable of reaching velocities as high as the speed of light, in other words, they are atoms, that lack electrons and contain high amounts of energy. Primary cosmic radiation is composed of protons, $\alpha$-particles (two protons and two neutrons), $\ {}^{-4}$e and heavy nuclei.
Cosmic rays enter to the Earth’s magnetosphere and other layers from all directions (primary cosmic rays). The primary CRs collide with particles in the Earth’s atmosphere, these violent collisions cause particles to break down into smaller particles (secondary cosmic rays). This process is repeated multiple times until the particles lose sufficient energy to become elemental particles or are unable to break down due to energy loss. This phenomenon is known as an air shower (Auger et al., 1939).
High-energy cosmic rays are produced by stochastic collisions within a magnetically turbulent plasma such as a supernova, and there are also low-energy cosmic rays from similar stars to our Sun (Sekido & Elliot, 1985).
Cosmic rays originating from sources beyond our solar system are detected on Earth’s surface, and their detection is affected by solar activity, which is reliant on the solar cycles. Furthermore, the Sun also generates cosmic rays within its corona.
The semblance has been used to compute the desired model. Semblance is a geophysical tool mainly used in exploration seismology and it can be defined as a correlation in Fourier terms or other transform between two traces, that is, semblance is a measure to compute the similarity between two traces.
@@ 2 Transport equation
The used solutions of transport equation in this document is based in Caballero-Lopez and Moraal (Caballero-Lopez & Moraal, 2004) article.
@@
The equation 1 is in terms of the omnidirectional dis- tribution function $f(r;P;t)$, $p$ is the particle’s moment, $t$ is the time, $r$ is the radial distance, $Q$ is the source of cosmic rays and $V$ the solar wind speed. $K$ is the diffusion tensor, that contains a parallel component $k_{\parallel}$, perpendicular $k_{\perp}$ and transverse $k_{T}$ and they describe the drifts. In the equation 2, $C$ (from the equation 3) is known as the Compton-Getting factor (Gleeson & Axford, 1968). This factor was introduced into the transport equation, because the flow, which is isotropic in the reference system at rest and is not the same in the solar wind system.
@@
Substituting 3 and 4 in 1 two possible solutions could be gotten:
@@ 3 Force Field
In the equation 5 the term $a$ corresponds to the convection, $b$ to the diffusion and drifts of the cosmic rays and $c$ are the changes of energy. To solve the previous equation, different approximations have been considered. If steady state, without sources and neglect the energy losses is assumed, from the equation 5 can be obtained:
$CVf-K\cdot\nabla f=cte=0$ (6)
For the equation 6 is valid, the term of adiabatic losses of the equation 7 is required to be smaller than the terms convective and diffusive:
@@
  • $(1/f)\partial f/\partial r<<C/r$ valid for small $r$, this means, in the internal heliosphere.
@@
The force field equation can be rewrite in terms of rigidity knowing that $P=cp/q$:
@@
The equation 11 implies that $f(r,p)$ is constant, equal to its value at the boundary, along an outline with characteristic equation $dP/dr=VP/3k$, in the space $(r,P)$. In the second term of the equation 8, the quantity $VP/k$ has units of potential per unit length, that is, field units, that is why it is called a force field solution. Considering the diffusion coefficient with the form:
@@
where $\phi$ is the force field parameter. When $k2\propto P$ and $\beta\approx 1$, the solution is reduced to the most commonly form utilized:
@@ 4 Convection-Diffusion
From the equation 5, $Q$ is not the adiabatic loss in a rest system:
@@
Where $M$ does not have unites.
Both solutions are introduced into the atmospheric yield solution, which is necessary to compute the modulation factor. However LIS is also necessary in order to compute the atmospheric yield solution.
@@ 5 Local Interstellar Spectral
The Local Interstellar Spectrum is a mathematical model created from observations or simulations that describe the arriving of the Cosmic Rays at outside the Heliosphere in energy terms.
The used LIS in these analyses were Lagner, Potgieter & Webber LIS in 2003, Burguer & Potgieter LIS in 2000, Garcia-Munoz, Mason & Simpson LIS in 1975 and Ghelfi, Barao, Derome & Maurin LIS in 2017.
@@ 5.1 Lagner, Potgieter & Webber LIS in 2003
This LIS is based on analytic solution of the Parker equation (Parker, 1965), in Voyager 1, Voyager 2 and Pioneer data at 70 Astronomical Units (AU) close to the Terminal Shock (TS) during minimum solar between 1987 and 1997.
The modulation developed by Parker included drift effects, that are related to the polarity of solar magnetic field:
@@
Where $j_{source}$ is the local source. Into the equation was introduced parallel, perpendicular, and asymmetric drift coefficients to describe the gradient and derivates of curvature in large-scale Heliosphere Magnetic Field (HMF).
The local sources are not considered and dependent time into spherical coordinates system like a combination model of diffusive shock acceleration and drift modulation in two spatial dimensions, neglecting any azimuthal dependence and contemplating that is symmetric around the equatorial field (Langner et al., 2003). The model offers several CR intensities that are interpreted as particles made of different composition. When the modulation is close or outside from TS is computed in rigidity terms. The protons model is the following:
@@ 5.2 Burguer & Potgieter LIS in 2000
The LIS was compared to ULYSSES data from September 1994 to July 1995 with a Fast Scanning of Latitude (FSL). The Parker equation was analytically solved considering an omnidirectional distribution (two dimensions), the source is located 100 UA, the Solar Wind’s (SW) velocity is $400\ \frac{km}{s}$ in equatorial plane and is increasing to $800\frac{km}{s}$ in polar zones. The HMF angle sheet is $15{{}^{\circ}}$, which is a good value to fit for solar modulation. The model was computed by Bibber (Bieber, 1999):
@@ 5.3 Garcia-Munoz, Mason and Simpson LIS in 1975
To compute this LIS was consulted IMP-5, IMP-7, and IMP-8 data, because the satellites can capture several particles and their energy rage is wider, then the select data were electron > $100MeV$ with quiet-time measurements.
The Parker’s equation was analytically solved considering symmetric spherical coordinates system in stationary state and drawing on Fisk method (Fisk et al., 1974). The LIS is the following:
@@
Where $j$ is particle flux, $T$ is kinetic energy and parameters are showing in the table 1:
@@ 5.4 Ghelfi, Barao, Derome and Maurin LIS in 2017
Used data were AMS-01 and AMS-02, BESS-Polar and PAMELA, because they were recent (in that epoch), but Voyager 1 data was also considered.
First, the author purposes the simplest modulation model to link unmodulated (IS) to modulated (TOA) qualities, that is a force field approximation Ghelfi et al. (2017):
@@
An analysis $\chi^{2}$ was necessary to fix the TOA flux with all the species $N_{j}(i)$ in this $t_{j}$, over all possible energy $E_{k}(i,\ j)$ and is giving by:
@@
Where IS parameters are free. For the previous mentioned was used power laws in total energy (O’Neill, 2006) or rigidity (Shikaze et al., 2007). TOA data for H and He are old and have more inconsistencies than new data, thus has applied a $\chi^{2}$ analysis to minimize errors, therefore was used package MINUIT minimisation package (James, 1998) from Root CERN libraries Sartini et al. (2010). Markov Chain Monte Carlo analysis (MCMC) was implemented from the GreAT package (Putze & Derome, 2014) to determine the correlation between credible intervals for $\phi$ and IS flux. MCMC analysis is based on Bayes theorem. Finally, all the obtained data is represented as an interpolated logarithmic-polynomial function, which is shown in the equation 28.
@@
In this paper were just applied four LIS, however more models were also considered, and it is worthwhile to mention them.
@@ 5.5 Boschini, Della Torre, Gervasi and more LIS in 2018
For the calculation of this LIS was used GALPROP and HeldMod to simulate CR data, but real data was required, so Voyager 1, BESS, PAMELA, AMS-01 and AMS-02 data were consulted.
GALPROP code use astronomical, particle physics and nuclear information to predict CR flux, X-ray, synchrotron emissions and their polarisation (Strong et al., 2007). Heliosphere propagation was computed by GALPROP and thanks to MCMC analysis, data was more realistic, after the information was introduced in HelMod program to obtain modulated data.
Previous data was compared with real information from the satellites with different energy content. The used information was a combination of Voyager 1, AMS-02, CREAM-I and ATIC-02, but CREAM-I and AMS-02 had better adjust because they offer minimal error. To compute the analytical LIS was used MCMC analysis from Eureqa [^2] http://www.nutonian.com/products/eureqa/ and the analytical model take the next form:
@@ 5.6 Vos & Potgieter LIS in 2015
The LIS was computed using observations from several PAMELA data during a solar minimum of 2006 to 2009. This LIS was computed from the Parker equation (Parker, 1965), which was solved as a ray distribution function in rigidity terms, time, position in 3D, in a heliocentric spherical coordinates where polar angles is equal to $90^{\circ}$. All short-term modulation effects are neglected (solar minimum conditions), average particle drift velocity is caused by gradients and curvature in Heliospheric Magnetic Field (HMF), there is a symmetric diffusion tensor and adiabatic energy changes except in heliosheath (Potgieter et al., 2012).
@@ 5.7 Moskalenko, Strong, Ormes and Potgieter LIS in 2002
Cosmic Ray data was simulated by DTUNUC Monte Carlo code (Ferrari et al., 1996) and data was recollected from two previous works (Tan & Ng, 1983) and (Tan & Ng, 1983). The Parker equation (Parker, 1965) was solved through Crank-Nicholson numerical solution (Crank & Nicolson, 1947) and accorded with the author a model more realistic was sought, then leaky box code was applied (weighted-slab). GALPROP was also used to compute 3D cosmic rays.
The Helium LIS was approximated to a Force Field, because of low energy approximation of the $\phi$ potential modulation has been selected using the CLIMAX neutron monitor.
@@
Finally, the LIS was subjected to a $\chi_{n}^{2}$ function to observe the quality of the adjusted data for each individual measurement (Moskalenko et al., 2002) is just shown in the equation 31:
@@
The LIS previously mentioned are shown in the figure 2 into the rigidity purpose parameters by each author:
@@ 6 Semblance
The semblance is the cross-relation between two traces, the computed value is interpreted as the similarity of two points contained in the signals that lie in the Fourier domain, however in this analysis was used Continuous Wavelet Transform (CWT) that represents a wave in two dimensions and measure the similarity between two signals.
@@
Where $s$ is scale, $u$ is displacement, $\Psi$ is the mother wavelet , and $*{}^{*}$ means complex conjugate and $t$ is the coordinate in time (Cooper & Cowan, 2008) the equation 32 can also be expressed in spatial terms (Teolis, 2017):
@@
The conventional semblance is a normalized coherence measure, first computed in 1969 (Taner & Koehler, 1969) A time later, Neidell and Taner (Neidell & Taner, 1971) found, that the coherence and semblance were two different operators, so semblance was expressed as shown in the equation 34:
@@
To reduce and make more smoothly the decays often is applied a boxcar filter, that can be rewritten as (Luo et al., 2015):
@@
There is an alternative expression for the conventional semblance, that can be expressed as a normalized correlation coefficient, but first a trace $r[j]$ should be defined as a summation over trace number:
@@
Now, the three equations 37, 38 and 39 can be rewritten as a conventional semblance (Luo et al., 2015):

### v3

@@ Abstract
We propose a quantitative model using the semblance method to evaluate differences in transport equation solutions and Local Interstellar Spectra (LIS) for Cosmic Ray (CR) modulation. The semblance method quantitatively compares transport solutions—convection-diffusion and force field—and LIS, providing insights into CR behavior.
By integrating these solutions into the atmospheric yield function (Caballero-Lopez & Moraal 2012), the model identifies a correlation between the modulation factor and sunspots, and an anti-correlation with the mean magnetic field. The force field solution shows stronger semblance values, aligning better with neutron detector data. Among LIS, models by Lagner, Potgieter & Webber (2003), and Ghelfi et al. (2017) are most consistent, while the Garcia-Munoz et al. (1975) LIS reveals notable variations.
This study underscores the semblance method as a critical tool for comparing CR modulation models, advancing predictive models of cosmic ray behavior influenced by solar and interstellar conditions.
Cosmic Rays — Convection-Diffusion solution — Force field solution — Semblance — Geophysics — Local interstellar spectral — Atmospheric yield function
@@ 1 Introduction
Primary cosmic rays are electrically charged particles and sometimes can reach velocities as high as the speed of light. Primary cosmic radiation is composed of protons, $\alpha$-particles (two protons and two neutrons), $\ {}^{-4}$e and heavy nuclei.
Primary cosmic rays enter the Earth’s magnetosphere and interact with the particles in the Earth’s atmosphere, those violent collisions produce smaller particles (secondary cosmic rays). This process is repeated multiple times until the particles lose sufficient energy to become elemental particles or are unable to break down due to energy loss. This phenomenon is known as an air shower (Auger et al., 1939).
High-energy cosmic rays are produced by stochastic collisions within a magnetically turbulent plasma such as a supernova, however there are also low-energy cosmic rays from similar stars to our Sun (Sekido & Elliot, 1985).
Cosmic rays originating from sources beyond our solar system are detected on Earth’s surface, and their detection is affected by solar activity, which is reliant on the solar cycles. Furthermore, the Sun also generates cosmic rays within its atmosphere.
The semblance method has been used to compute the desired model. Semblance is a geophysical tool mainly used in exploration seismology, and it can be defined as a correlation in Fourier terms or other transforms between two traces, that is, semblance is a measure to compute the similarity between two traces.
@@ 2 Transport equation
The solutions of the transport equation used in this document are based on Caballero-Lopez and Moraal (Caballero-Lopez & Moraal, 2004) article.
@@
The equation 1 is in terms of the omnidirectional distribution function $f(r;P;t)$, where $p$ is the moment of the particle, $t$ is the time, $r$ is the radial distance, $Q$ is the source of cosmic rays and $V$ the solar wind speed. $K$ is the diffusion tensor, that contains a parallel component $k_{\parallel}$, perpendicular $k_{\perp}$ and transverse $k_{T}$, and they describe the drifts. In the equation 2, $C$ (from the equation 3) is known as the Compton-Getting factor (Gleeson & Axford, 1968). This factor was introduced into the transport equation, because the flow is isotropic in the reference system at rest and is not the same in the solar wind system.
@@
Substituting 3 and 4 into 1 two possible solutions could be obtained:
@@ 3 Force Field
In equation 5 the term $a$ corresponds to the convection, $b$ to the diffusion and drifts of the cosmic rays, and $c$ to the changes in energy. To solve the previous equation, different conditions have been considered. If a steady state, no sources, and negligible energy losses are assumed, the following can be obtained from equation 5:
$CVf-K\cdot\nabla f=cnst.=0$ (6)
Equation 6 is valid when the term for adiabatic losses in equation 7 is smaller than the terms convective and diffusive:
@@
  • $(1/f)\partial f/\partial r<<C/r$ valid for small $r$, that is, in the internal heliosphere.
@@
The force field equation in terms of rigidity and knowing that $P=cp/q$:
@@
Equation 11 implies that $f(r,p)$ is constant, equal to its value at the boundary, along a contour with characteristic equation $dP/dr=VP/3k$, in the space $(r,P)$. In the second term of the equation 8, the quantity $VP/k$ has units of potential per unit length, that is, force units, this is why it is called a force field solution. Considering the diffusion coefficient with the form:
@@
where $\phi$ is the force field parameter. When $k2\propto P$ and $\beta\approx 1$, the solution is reduced to the most commonly used form:
@@ 4 Convection-Diffusion
From equation 5, $Q$ is not the adiabatic loss in a rest system:
@@
Where $M$ does not have units.
Both solutions are introduced into the atmospheric yield solution, which is necessary to compute the modulation factor. However, LIS is also necessary in order to compute the atmospheric yield solution.
@@ 5 Local Interstellar Spectral
The Local Interstellar Spectrum is a mathematical model created from observations or simulations and describe the arrival of the Cosmic Rays outside the Heliosphere in energy terms.
The LIS models used in these analyses were Lagner, Potgieter & Webber LIS in 2003, Burguer & Potgieter LIS in 2000, Garcia-Munoz, Mason & Simpson LIS in 1975 and Ghelfi, Barao, Derome & Maurin LIS in 2017.
@@ 5.1 Lagner, Potgieter & Webber LIS in 2003
This LIS is based on an analytic solution of the Parker equation (Parker, 1965), using Voyager 1, Voyager 2 and Pioneer data at 70 Astronomical Units (AU) close to the Terminal Shock (TS) during the solar minimum between 1987 and 1997.
The modulation developed by Parker included drift effects which are related to the polarity of solar magnetic field:
@@
Where $j_{source}$ is the local source. In the equation, parallel, perpendicular, and asymmetric drift coefficients were introduced to describe the gradients and drifts of curvature in large-scale Heliospheric Magnetic Field (HMF).
The local sources are not considered and are time-dependent in a spherical coordinate system modeled as a combination of diffusive shock acceleration and drift modulation in two spatial dimensions, neglecting any azimuthal dependence and assuming a symmetrical equatorial field (Langner et al., 2003). The model offers several CR intensities, which are interpreted as particles of different compositions. When the modulation is computed close to or outside the TS, it shall be computed in rigidity terms. The protons model is the following:
@@ 5.2 Burguer & Potgieter LIS in 2000
The LIS was compared to ULYSSES data from September 1994 to July 1995 with a Fast Scanning of Latitude (FSL). The Parker equation was analytically solved considering an omnidirectional distribution (two dimensions), the source is located 100 UA, the velocity of the Solar Wind (SW) is $400\ \frac{km}{s}$ in equatorial plane, and it is increasing to $800\frac{km}{s}$ in polar zones. The HMF angle sheet is $15{{}^{\circ}}$, which is a good value to fit for solar modulation. The model was computed by Bibber (Bieber, 1999):
@@ 5.3 Garcia-Munoz, Mason and Simpson LIS in 1975
To compute this LIS, IMP-5, IMP-7, and IMP-8 data were consulted, because the satellites can capture several particles and their energy range is the widest. The selected data” were electrons > $100MeV$ with quiet-time measurements.
The Parker’s equation was analytically solved considering a symmetric spherical coordinate system in steady state and drawing on Fisk’s method (Fisk et al., 1974). The LIS is the following:
@@
Where $j$ is particle flux, $T$ is kinetic energy and parameters are shown in Table 1:
@@ 5.4 Ghelfi, Barao, Derome and Maurin LIS in 2017
The data used were AMS-01 and AMS-02, BESS-Polar and PAMELA, because they were recent (at that time), but Voyager 1 data was also considered.
First, the author proposes the simplest modulation model to link unmodulated (IS) to modulated (TOA) quantities, which is a force field approximation Ghelfi et al. (2017):
@@
An analysis $\chi^{2}$ was necessary to fix the TOA flux with all the species $N_{j}(i)$ in this $t_{j}$, over all possible energy $E_{k}(i,\ j)$ and is given by:
@@
Where IS parameters are free. For the previously mentioned cases, power laws in total energy (O’Neill, 2006) or rigidity (Shikaze et al., 2007). TOA data for H and He are old and have more inconsistencies than new data, thus, a $\chi^{2}$ analysis was applied to minimize errors. Therefore, the MINUIT minimization package was used (James, 1998) from Root CERN libraries Sartini et al. (2010). Markov Chain Monte Carlo analysis (MCMC) was implemented from the GreAT package (Putze & Derome, 2014) to determine the correlation between credible intervals for $\phi$ and IS flux. MCMC analysis is based on Bayes’ theorem. Finally, all the obtained data is represented as an interpolation of logarithmic-polynomial functions, as shown in equation 28.
@@
In this paper, just four LIS were applied, however more models were also considered, and it is worthwhile to mention them.
@@ 5.5 Boschini, Della Torre, Gervasi and more LIS in 2018
For the calculation of this LIS the GALPROP and HelMod programs were used to simulate CR data, but real data was required, so Voyager 1, BESS, PAMELA, AMS-01 and AMS-02 data were consulted.
The GALPROP code uses astronomical, particle physics and nuclear information to predict CR flux, X-ray, synchrotron emissions and their polarization (Strong et al., 2007). Heliosphere propagation was computed by GALPROP and thanks to the MCMC analysis, the data became more realistic. Afterward, the information was introduced into the HelMod program to obtain modulated data.
Previous data was compared with real information from the satellites with different energy content. The used information was a combination of Voyager 1, AMS-02, CREAM-I and ATIC-02, but CREAM-I and AMS-02, because they had better adjustments and offered minimal errors. To compute the analytical LIS, MCMC analysis from Eureqa [^2] http://www.nutonian.com/products/eureqa/ was used, the outcome is given:
@@ 5.6 Vos & Potgieter LIS in 2015
The LIS was computed using observations from PAMELA data during a solar minimum from 2006 to 2009. This LIS was computed from the Parker equation (Parker, 1965), which was solved as a ray distribution function in rigidity terms, time, position in 3D, in heliocentric spherical coordinates where the polar angle is equal to $90^{\circ}$. All short-term modulation effects were neglected (solar minimum conditions). The average particle drift velocity is caused by gradients and curvature in Heliospheric Magnetic Field (HMF), there is a symmetric diffusion tensor and adiabatic energy changes except in the heliosheath (Potgieter et al., 2012).
@@ 5.7 Moskalenko, Strong, Ormes and Potgieter LIS in 2002
The cosmic ray data was simulated using the DTUNUC Monte Carlo code (Ferrari et al., 1996) , and additional data was collected from two previous works (Tan & Ng, 1983). The Parker equation (Parker, 1965) was solved through Crank-Nicholson numerical solution (Crank & Nicolson, 1947) and according to the author: a more realistic model was sought. Thus, the leaky box code (weighted-slab) was applied. GALPROP was also used to compute 3D cosmic rays.
The Helium LIS was approximated using a force field solution due to the low-energy approximation, and the $\phi$ potential modulation was selected using the CLIMAX neutron monitor.
@@
Finally, the LIS was subjected to a $\chi_{n}^{2}$ function to observe the quality of the adjusted data for each individual measurement (Moskalenko et al., 2002) as shown in equation 31:
@@
The previously mentioned LIS are shown in the figure 2 using the rigidity-purpose parameters defined by each author:
@@ 6 Semblance
The semblance is the cross-relation between two traces, the computed value is interpreted as the similarity of two points contained in the traces within the Fourier domain. However, in this analysis, the Continuous Wavelet Transform (CWT) was used which represents a wave in two dimensions: data number and wavelength.
@@
Where $s$ is the scale, $u$ is the displacement, $\Psi$ is the mother wavelet , and means complex conjugate and $t$ is the coordinate in time (Cooper & Cowan, 2008) Equation 32 can also be expressed in spatial terms (Teolis, 2017):
@@
The conventional semblance is a normalized coherence measure, first computed in 1969 (Taner & Koehler, 1969),later, Neidell and Taner (Neidell & Taner, 1971) found that the coherence and semblance were two different operators. Thus, semblance was expressed as shown in the equation 34:
@@
To reduce and smooth the decays a boxcar filter is often applied, that can be written as (Luo et al., 2015):
@@
There is an alternative expression for the conventional semblance, which can be expressed as a normalized correlation coefficient. However, first, a trace $r[j]$ should be defined as a summation over the trace number:
@@
Now, the three equations 37, 38 and 39 can be expressed as a conventional semblance (Luo et al., 2015):